• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E. Cisbani

    Log entry time 08:06:11 on November 7,2008

    Entry number 247294

    keyword=Owl Shift Summary

    Owl Shift Summary

    00:30 BeO centered in Tune Beam

    00:40 Asked for Harp Scan at 5 uA

    00:15 Get 5 uA on Holey Carbon target (raster on start 3x3-MCC)

    02:00 No hole in the raster plot, and lots of warning about number of fired channels

    02:15 Vince paged, try to undestand the raster problem

    02:30 Alex paged, he will come in 5-10 minutes, in the manwhile, we start the detector checkout

    04:30 MCC: Tune beam position

    04:45 Start RICH tests

    06:45 Raster study on BeO (raster on then off)
    problem with ROC9; current decreased with raster off

    07:10 Strangely the ABU is 0, while we got data (for commissioning)